OpenStudy (anonymous):

The endpoints of segment GE are located at G(–6, –4) and E(4, 8). Using slope-intercept form, write the equation of line GE

OpenStudy (anonymous):

none

OpenStudy (anonymous):

y=-6/5x+64/5 y=5/6x+14/3 y=6/5x+16/5 y=6/5x+64/5

OpenStudy (anonymous):

ye

OpenStudy (elson):

find the slope by delta y over delta x. The slope is 12/10=6/5. Then plug in one of the point into y=6/5x+c. C is 16/5. The third choice is the answer

OpenStudy (elson):

Just helping people by teaching them, not direct answers

OpenStudy (skullpatrol):

you told the asker the correct answer choice?

OpenStudy (anonymous):

flutter them elson, you the real MVP

OpenStudy (anonymous):

fck*

OpenStudy (anonymous):

even tho i guessed on my entire test a failed it with a 20 lol

OpenStudy (elson):

Glad to be of help

OpenStudy (skullpatrol):

Nobody can help you @sebastian_starnes if you don't want to try